# STATE ToxiCN: A Benchmark for Span-level Target-Aware Toxicity Extraction in Chinese Hate Speech Detection

### v1

@@ Abstract
The proliferation of hate speech has caused significant harm to society. The intensity and directionality of hate are closely tied to the target and argument it is associated with. However, research on hate speech detection in Chinese has lagged behind, and existing datasets lack span-level fine-grained annotations. Furthermore, the lack of research on Chinese hateful slang poses a significant challenge. In this paper, we provide a solution for fine-grained detection of Chinese hate speech. First, we construct a dataset containing Target-Argument-Hateful-Group quadruples (STATE ToxiCN), which is the first span-level Chinese hate speech dataset. Secondly, we evaluate the span-level hate speech detection performance of existing models using STATE ToxiCN. Finally, we conduct the first study on Chinese hateful slang and evaluate the ability of LLMs to detect such expressions. Our work contributes valuable resources and insights to advance span-level hate speech detection in Chinese. [^1] Code and datasets are publicly available at https://github.com/shenmeyemeifashengguo/STATE-ToxiCN.
@@ 1 Introduction
With the popularity of social media, user-generated content has experienced explosive growth, and hate speech has also flourished. Hate speech refers to harmful statements that express hatred or incite harm against specific groups or individuals based on race, religion, gender, region, sexual orientation, physical characteristics, and other factors Bilewicz and Soral (2020). These statements are often driven by group hysteria, personal experiences, and social trends, and are characterized by compulsiveness, bullying, and incitement. Approximately 941 million people, or 12% of the global population, speak Mandarin Chinese as their first language Eberhard et al. (2024). The issue of effectively detecting and managing Chinese hate speech has become a topic of widespread concern.
The intensity and directionality of hate speech are closely tied to the target and argument it is associated with Cowan and Hodge (1996). In recent years, research on hate speech detection in English has gradually shifted from post-level Davidson et al. (2017); Founta et al. (2018) to span-level Pavlopoulos et al. (2021); Mathew et al. (2021); Zampieri et al. (2023). However, in the Chinese domain, span-level hate speech detection remains unexplored, significantly limiting the development of Chinese hate speech detection. This task faces unique challenges within the Chinese linguistic environment. Firstly, Chinese syntax’s flexible word order makes identifying the target and argument ambiguous. For example, as shown in Exp.1 in Table 1, the use of inversion in Chinese disrupts the conventional subject-verb-object structure. Secondly, the absence of word delimiters in Chinese, such as the spaces used in English, complicates word segmentation. Furthermore, the scarcity of annotated resources exacerbates this issue, highlighting the urgent need for research in span-level Chinese hate speech detection.
@@
The ideographic nature of Chinese, with its abundance of synonyms and near-synonyms, leads to highly varied hateful slang. Chinese hateful slang often evades model detection through various techniques such as homophonic substitutions, character splitting and merging, and historical allusions, which severely compromise the accuracy of existing models. For example, in Exp.3 from Table 1, the target uses a merge-word technique to evade detection. However, previous work has only provided words, lacking interpretable annotations. To address this gap, we collect and summarize commonly used hateful slang from real-world online forums, providing detailed annotations to construct the first annotated lexicon for Chinese hateful slang. This lexicon offers a reliable benchmark for Chinese hate speech detection and lays a foundation for improving model performance.
To address the lack of resources for span-level Chinese hate speech research, we introduce a Span-level Target-Aware Toxicity Extraction dataset (STATE ToxiCN), a novel dataset containing 8029 posts and 9533 quadruples addressing sexism, racism, regional bias, and anti-LGBTQ sentiments. Using this dataset, we evaluate the performance of LLMs in span-level Chinese hate speech detection. Specially, we annotate: 1) extraction of the targets and arguments from the post, 2) determination of whether each Target-Argument pair constitutes hate speech, and 3) classification of the groups for hateful Target-Argument pairs.
Moreover, we systematically summarize and annotate commonly used hateful slang from real-world online forums to address the challenges of identifying hateful slang. We compile a comprehensive annotated lexicon, labeling each hateful slang with its frequent groups and providing detailed explanations of its usage and contextual meaning. This resource not only helps to understand how hateful slang are used to disguise hate speech but also provides valuable annotated data for evaluating and improving the ability of LLMs to detect Chinese hateful slang. The main contributions of this work are summarized as follows:
@@
  • We construct an annotated lexicon of commonly used hateful slang in online forums, including their frequent groups and contextual explanations as annotations.
@@ 2 Related Work
Hate speech detection is a critical task in NLP that has attracted considerable attention recently. Researchers have increasingly turned to pre-trained models to address this issue Caselli et al. (2020); Hanu and Unitary team (2020); Zhou et al. (2021); AlKhamissi et al. (2022); Ali et al. (2022). To facilitate progress in this field, several datasets tailored to hate speech detection have been developed Waseem and Hovy (2016); Davidson et al. (2017); Founta et al. (2018); Hartvigsen et al. (2022). Pavlopoulos et al. (2021) introduced span-level hate speech detection, while the TBO dataset advanced the field by pioneering the extraction of Target-Argument-Harmful triples Zampieri et al. (2023). However, Chinese hate speech detection remains significantly underdeveloped.
@@
While some Chinese hate speech datasets exist, these efforts remain limited to the post-level. TOCP and TOCAB, derived from Taiwan’s PTT platform, focus on detecting profanity and abusive language Chung and Lin (2021). The Sina Weibo Sexism Review (SWSR) centers on identifying sexism Jiang et al. (2022). The Chinese Offensive Language Dataset (COLD) categorizes sentences into types such as individual attacks and anti-bias Deng et al. (2022). Zhou et al. (2022) introduce CDial-Bias, the first annotated dataset specifically addressing social bias in Chinese dialogues. Lu et al. (2023) present ToxiCN, a dataset encompassing both explicit and implicit toxic language samples.
While previous works offer quality corpora, span-level Chinese hate speech detection remains unexplored. STATE ToxiCN is the first such dataset, annotating Target-Argument-Hateful- Group quadruples within posts. Unlike previous lexicons, we provide interpretative annotations and targeted group labels, creating a comprehensive lexicon of Chinese hateful slang.
@@ 3.3.1 Annotation Guidelines
During the annotation process, we adhere to strict guidelines and implement multi-stage quality control to ensure the accuracy and consistency of the annotations, ultimately constructing a high-quality span-level Chinese hate speech dataset. First, we develop detailed annotation guidelines, including standards for extracting targets and arguments (Target-Argument Pair), criteria for determining hatefulness (Hateful), and methods for classifying groups (Group):
@@
A combination of the target and the argument extracted from the post. A single post may contain more than one Target-Argument Pair.
@@
During the construction of the annotated lexicon for Chinese hateful slang, we follow well-defined guidelines and employ systematic quality control to ensure reliability. This process include detailed criteria for identifying, categorizing, and annotating hateful slang terms, with a focus on their contextual usage and targeted implications:
@@ 3.3.3 Annotation Procedure
Each text is independently annotated by at least two annotators, who followed a unified annotation guideline to extract Target-Argument pairs, determine the hatefulness, and classify the groups. After the initial annotation phase, a portion of the samples is regularly selected for cross-validation, during which other annotators re-annotate these samples to ensure consistent understanding of the rules and standards across annotators. This approach allows us to identify and resolve potential biases or discrepancies in the annotations in a timely manner.
For disputed samples, where significant differences in annotation results are observed between annotators, the samples are submitted to an arbitration team composed of domain experts for review. The expert team analyze the disputed samples in depth, taking into account the textual context and the annotation guidelines, and determined the most reasonable annotation results. After annotation, we evaluate inter-annotator agreement (IAA) on STATE ToxiCN, with Fleiss’ kappa scores for each hierarchy detailed in Table 3.
@@
Additionally, the document serves as a knowledge-sharing platform, providing the annotation team with consistent references and concrete examples for handling complex or ambiguous posts. This mechanism improves annotation efficiency and enhances the consistency of the annotations. The dynamically updated lexicon of annotated Chinese hateful slang become a valuable resource for research on Chinese hate speech detection.
@@ 3.4 Data Description
STATE ToxiCN dataset contains a total of 8,029 annotated posts, among which 4,942 posts include hateful content, accounting for 61.55%. A total of 9,533 quadruples are annotated, with 6,034 of them containing hateful information, making up 63.60%. We present the statistical details of STATE ToxiCN in Table 4. Gender, Region, and Race are the three most common group types in the dataset. Additionally, "multi-group" refers to Target-Argument pairs that involve hatred directed at multiple target groups, with a total of 854 such instances, accounting for 8.96%. In addition, the annotated lexicon includes 330 Chinese hateful slang terms collected from real online forums. More annotated examples can be found in Appendix A.
@@ 4.1 Baselines
Open-source Models: mT5-base Xue (2020), Mistral-7B Jiang et al. (2023), LLaMA3-8B AI@Meta (2024), Qwen2.5-7B Team (2024); Safety-domain Models: ShieldLM-14B-Qwen Zhang et al. (2024), ShieldGemma-9B Zeng et al. (2024), and Closed-source LLMs: LLaMA3-70B AI@Meta (2024), Qwen2.5-72B Team (2024), Gemini-1.5-Pro Team et al. (2023), Claude-3.5-Sonnet Anthropic (2024), GPT-4o OpenAI (2024), DeepSeek-v3 DeepSeek-AI et al. (2024).
@@
We employ the Macro-F1 scores as the main evaluation metrics. The lack of word delimiters in Chinese introduces ambiguity in the extraction of targets and arguments. To obtain more accurate evaluation results, we therefore utilize both hard and soft matching metrics as our .
@@ 4.3 Experiment Settings
We conduct fine-tuning on open-source and safety-domain models, with training and testing set sizes detailed in Table 5. The fine-tuning process was performed using only a basic prompt, which included task definitions, output formats, and specific prediction requirements for all elements. We evaluate the performance of closed-source LLMs by calling their APIs. In addition to the aforementioned basic prompt, we also provided a hate speech example and a non-hate speech example. Further details regarding these requirements are provided in Appendix C.
@@ 5.3.1 Impact of Hateful Slang on Chinese Hate Speech Detection
To evaluate the impact of hate slang on span-level Chinese hate speech detection models, we select 502 posts containing hate slurs from the test set for focused testing. By comparing the model performance on this subset of data with its performance on the full test set, we can gain a clearer understanding of how hate slurs affect model performance. The experimental results are presented in Table 7. The superscript numbers in the table indicate the difference in F1 score between the model’s performance on the selected subset and its performance on the full test set. Green indicates a decrease in F1 score on the subset compared to the full dataset, while red indicates an increase.
@@
Surprisingly, LLM APIs demonstrate superior performance on the subset of the test set containing hateful slang. With the exception of a slight decrease in the soft matching metric for argument identification, nearly all other metrics show an improvement. We believe that this is primarily due to two factors: First, hateful slang often appears as the target, and the explicit target terms effectively compensate for the limitations of LLM APIs in span boundary recognition. Second, the identification of hateful slang often requires rich background knowledge; for example, Example 2 in the figure requires knowledge of Chinese history and related common usages, which is precisely where LLM APIs excel. Therefore, hateful slang, which is difficult for fine-tuned models to leverage, becomes an effective source of information for LLM APIs.
@@ 5.3.2 Understanding Capabilities of LLMs on Chinese Hateful Slang
To investigate the understanding capabilities of large language models regarding Chinese hateful slang, we conducted experiments using two relatively high-performing models, ShieldGemma-9B and DeepSeek-v3. In these experiments, we provided the models with hateful terms and their surrounding sentence context, asking them to generate explanations of the hateful slang and identify the potentially affected groups. The experimental results are detailed in Figure 1, with specific experimental details provided in Appendix B.
@@ 6 Conclusions
Different from previous work focusing on template design, we address implicit sentiment detection from a human cognitive perspective. In this paper, we propose a straightforward and effective strategy-level approach inspired by dual-process theory, STATE ToxiCN, which detects implicit sentiment by a dual-system reasoning framework in the ASQP task. Firstly, we employ commonsense reasoning tools to estimate human reactions and form intuitive reaction augmentation samples. Secondly, we integrate the augmented text with analytic system-based instruction and use this information to enable the model to further integrate contextual understanding and intuitive reaction augmentation samples, resulting in a comprehensive construction of the sentiment quads. The experimental results show that our STATE ToxiCN can achieve state-of-the-art performance compared to existing baseline methods. We also conduct experiments in various scenarios to thoroughly demonstrate the effectiveness, universality, and robustness of STATE ToxiCN.

### v2

@@ Abstract
The proliferation of hate speech has caused significant harm to society. The intensity and directionality of hate are closely tied to the target and argument it is associated with. However, research on hate speech detection in Chinese has lagged behind, and existing datasets lack span-level fine-grained annotations. Furthermore, the lack of research on Chinese hateful slang poses a significant challenge. In this paper, we provide two valuable fine-grained Chinese hate speech detection research resources. First, we construct a Span-level Target-Aware Toxicity Extraction dataset (STATE ToxiCN), which is the first span-level Chinese hate speech dataset. Secondly, we evaluate the span-level hate speech detection performance of existing models using STATE ToxiCN. Finally, we conduct the first study on Chinese hateful slang and evaluate the ability of LLMs to understand hate semantics. Our work contributes valuable resources and insights to advance span-level hate speech detection in Chinese.  [^1] Code and datasets are publicly available at https://github.com/shenmeyemeifashengguo/STATE-ToxiCN.
@@ 1 Introduction
With the popularity of social media, user-generated content has experienced explosive growth, and hate speech has also flourished. Hate speech refers to harmful statements that express hatred or incite harm against specific groups or individuals based on race, religion, gender, region, sexual orientation, physical characteristics, and other factors Bilewicz and Soral (2020). Due to its damaging impact on individuals and society, hate language is now widely considered as a problem of increasing importance Silva et al. (2016). Recently, researchers have been actively engaged in the study of hate speech detection, and this research has gradually shifted from post-level Ahn et al. (2024); AlKhamissi et al. (2022b) to span-level Pavlopoulos et al. (2021); Mathew et al. (2021); Zampieri et al. (2023).
Approximately 941 million people, or 12% of the global population, speak Mandarin Chinese as their first language Eberhard et al. (2024). However, research on Chinese hate speech detection lags significantly behind. There are two key issues that remain unresolved. Firstly, existing research Deng et al. (2022); Jiang et al. (2022); Zhou et al. (2022) is limited to the post-level, leaving span-level Chinese hate speech detection unexplored. The intensity and directionality of hate speech are closely tied to the target and argument it is associated with Cowan and Hodge (1996). In the Chinese linguistic context, challenges in span-level hate speech detection arise due to flexible word order and the absence of word delimiters like spaces. These issues complicate the identification of targets and arguments. For example, as shown in Exp.1 in Table 1, the use of inversion in Chinese disrupts the conventional subject-verb-object structure. To fill this gap, we construct a span-level Chinese hate speech dataset by annotating the Target-Argument-Hateful-Group quadruples in posts.
@@
Secondly, although Chinese hate lexicons have been developed Jiang et al. (2022); Lu et al. (2023), the absence of interpretable annotations hinders a deeper understanding of hate speech. As an ideographic language, Chinese has a wealth of synonyms and near-synonyms Mair (1991), making the forms and types of hateful slang diverse and difficult to capture. Chinese hateful slang often evades model detection through techniques such as homophonic substitution, character splitting and merging, and historical allusions Xiao et al. (2024). For example, in Exp. 3 from Table 1, the target uses the merging technique to evade detection. To address this issue, we collect commonly used hateful slang from real-world online forums, providing detailed annotations to construct the first annotated lexicon for Chinese hateful slang. This lexicon serves as a foundational resource for research into the understanding of Chinese hate semantics.
To address the lack of resources for span-level Chinese hate speech research, we introduce a Span-level Target-Aware Toxicity Extraction dataset (STATE ToxiCN), a novel dataset containing 8029 posts and 9533 Target-Argument-Hateful-Group quadruples addressing sexism, racism, regional bias, and anti-LGBTQ sentiments. Using this dataset, we evaluate the performance of LLMs in span-level Chinese hate speech detection. Specially, we annotate: 1) extraction of the targets and arguments from the post, 2) determination of whether each Target-Argument pair constitutes hate speech, and 3) classification of the groups for hateful Target-Argument pairs.
Moreover, we systematically summarize and annotate commonly used hateful slang from real-world online forums to address the challenges of identifying hateful slang. We compile a comprehensive annotated lexicon, labeling each hateful slang with its frequent groups and providing detailed explanations of its usage and contextual meaning. This is the first annotated lexicon of Chinese hateful slang with interpretative annotations. This resource not only helps to understand how hateful slang are used to disguise hate speech but also provides valuable annotated data for evaluating and improving the ability of LLMs to detect Chinese hateful slang. The main contributions of this work are summarized as follows:
@@
  • We construct an annotated lexicon of commonly used hateful slang in real-world online forums to evaluate LLMs’ capability in understanding Chinese hate speech semantics.
@@ 2 Related Work
Hate speech detection is a critical task in NLP that has attracted considerable attention recently. Researchers have increasingly turned to pre-trained models to address this issue Caselli et al. (2020); Hanu and Unitary team (2020); Zhou et al. (2021); AlKhamissi et al. (2022a); Ali et al. (2022). To facilitate progress in this field, several datasets tailored to hate speech detection have been developed Waseem and Hovy (2016); Davidson et al. (2017); Founta et al. (2018); Hartvigsen et al. (2022). Pavlopoulos et al. (2021) introduced span-level hate speech detection, while the TBO dataset advanced the field by pioneering the extraction of Target-Argument-Harmful triples Zampieri et al. (2023). However, Chinese hate speech detection remains significantly underdeveloped.
@@
While some Chinese hate speech datasets exist, these efforts remain limited to the post-level. TOCP and TOCAB, derived from Taiwan’s PTT platform, focus on detecting profanity and abusive language Chung and Lin (2021). The Sina Weibo Sexism Review (SWSR) centers on identifying sexism Jiang et al. (2022). The Chinese Offensive Language Dataset (COLD) categorizes sentences into types such as individual attacks and anti-bias Deng et al. (2022). Zhou et al. (2022) introduce CDial-Bias, the first annotated dataset specifically addressing social bias in Chinese dialogues. Lu et al. (2023) present ToxiCN, a dataset encompassing both explicit and implicit toxic language samples. Xiao et al. (2024) introduce a dataset to evaluate LLMs’ robustness against cloaking perturbations.
While previous works offer quality corpora, span-level research remains unexplored. STATE ToxiCN is the first such dataset, annotating Target-Argument-Hateful- Group quadruples. Unlike existing lexicons, we offer interpretative annotations and targeted group labels, creating a comprehensive Chinese hateful slang lexicon.
@@ 3.3.1 Annotation Guidelines
During the annotation process, we establish guidelines and implement multi-stage quality control to ensure the consistency of the annotations, aiming to label Target-Argument-Hateful-Group quadruples in the posts. First, we develop detailed annotation guidelines, including standards for extracting targets and arguments (Target-Argument Pair), criteria for determining hatefulness (Hateful), and methods for classifying groups (Group):
@@
A span consists of both the target and its corresponding argument extracted from the post. A single post may contain more than one Target-Argument Pair.
@@
For example, in Exp. 2 from Table 1, Target-Argument Pair is “男同(Gay people)” and “艾滋高发群体(a high-risk group for HIV)”. This Target-Argument Pair constitutes hate against the LGBTQ community and people living with AIDS, so the Hateful label is “hate” and the Group label is “LGBTQ, others”. The quadruple is [ 男同 | 艾滋高发群体 | hate | LGBTQ, others ].
Regarding the annotation of the Chinese hateful slang lexicon, we also establish annotation guidelines for identifying, categorizing, and labeling the terms, with a focus on their frequent groups (Group) and contextual explanations (Explanation):
@@ 3.3.3 Annotation Procedure
Each text is independently annotated by at least two annotators, who followed a unified annotation guideline to extract Target-Argument pairs, determine the hatefulness, and classify the groups. After the initial annotation phase, 20% of the samples are regularly selected for cross-validation, during which other annotators re-annotate these samples to ensure a consistent understanding of the rules and standards across annotators. This approach allows us to identify and resolve potential biases or discrepancies in the annotations in a timely manner.
For disputed samples with significant annotation differences, an arbitration team of domain experts reviews the samples, considering the textual context and annotation guidelines to determine the most appropriate annotation. We explore inter-annotator agreement (IAA) on STATE ToxiCN, with Fleiss’ kappa scores Fleiss (1971) for each hierarchy detailed in Table 3. The detailed analysis can be found in Appendix D.
@@
Additionally, the shared online document serves as a knowledge-sharing platform, providing the annotation team with consistent references and concrete examples for handling complex or ambiguous posts. This mechanism improves annotation efficiency and enhances the consistency of the annotations. This is the first Chinese hateful slang lexicon with interpretable annotations. This lexicon provides valuable research resources for span-level Chinese hate speech semantic understanding.
@@ 3.4 Data Description
STATE ToxiCN dataset contains a total of 8,029 annotated posts, among which 4,942 posts include hateful content, accounting for 61.55%. A total of 9,533 quadruples are annotated, with 6,034 of them containing hateful information, making up 63.60%. We present the statistical details of STATE ToxiCN in Table 4. Gender, Region, and Race are the three most common group types in the dataset. Additionally, "multi-group" refers to Target-Argument pairs that involve hatred directed at multiple target groups, with a total of 854 such instances, accounting for 8.96%. In addition, the annotated lexicon includes 830 Chinese hateful slang terms collected from real online forums. More annotated examples can be found in Appendix A.
@@ 4.1 Baselines
Open-source Models: mT5-base Xue (2020), Mistral-7B Jiang et al. (2023), LLaMA3-8B AI@Meta (2024), Qwen2.5-7B Team (2024); Safety-domain Models: ShieldLM-14B-Qwen Zhang et al. (2024), ShieldGemma-9B Zeng et al. (2024), and Closed-source LLMs: LLaMA3-70B AI@Meta (2024), Qwen2.5-72B Team (2024), Gemini-1.5-Pro Team et al. (2023), Claude-3.5-Sonnet Anthropic (2024), GPT-4o OpenAI (2024), DeepSeek-v3 Liu et al. (2024).
@@
Due to the ambiguity of Chinese span boundaries, a single evaluation metric may not accurately assess the performance of models in span-level Chinese hate speech detection. To obtain more accurate evaluation results, we therefore utilize both hard and soft matching metrics. We employ the Macro-F1 scores as the main evaluation metrics.
@@ 4.3 Experiment Settings
We conduct fine-tuning on open-source and safety-domain models, with training and testing set sizes detailed in Table 5. The fine-tuning process was performed using only a basic prompt, which included task definitions, output formats, and specific prediction requirements for all elements. We evaluate the performance of closed-source LLMs by calling their APIs. In addition to the aforementioned basic prompt, we also provided a hate speech example and a non-hate speech example. Further details regarding these requirements are provided in Appendix B.
@@ 5.3.1 Impact of Hateful Slang on Chinese Hate Speech Detection
To evaluate the impact of hate slang on span-level Chinese hate speech detection models, we remove the posts without hateful slang from the test set and obtain a subset of 502 posts containing hateful slang for focused testing. By comparing the model performance on this subset of data with its performance on the full test set, we can gain a clearer understanding of how hate slurs affect model performance. The experimental results are presented in Table 7. The superscript numbers in the table indicate the difference in F1 score between the model’s performance on the selected subset and its performance on the full test set. Green indicates a decrease in F1 score on the subset compared to the full dataset, while red indicates an increase.
@@
LLM APIs demonstrate superior performance on the subset of the test set containing hateful slang. With the exception of a slight decrease in the soft matching metric for argument identification, nearly all other metrics show an improvement. We believe that this is primarily due to two factors: First, hateful slang often appears as the target, and the explicit target terms effectively compensate for the limitations of LLM APIs in span boundary recognition. Second, the identification of hateful slang often requires rich background knowledge; for example, Example 2 in the figure requires knowledge of Chinese history and related common usages, which is precisely where LLM APIs excel. Therefore, hateful slang, which is difficult for fine-tuned models to leverage, becomes an effective source of information for LLM APIs.
@@ 5.3.2 Understanding Capabilities of LLMs on Chinese Hateful Slang
To investigate the understanding capabilities of large language models regarding Chinese hateful slang, we conducted experiments using two relatively high-performing models, ShieldGemma-9B and DeepSeek-v3. In these experiments, we provided the models with hateful terms and their surrounding sentence context, asking them to generate explanations of the hateful slang and identify the potentially affected groups. The experimental results are detailed in Figure 1, with specific experimental details provided in Appendix C.
@@ 6 Conclusions
With the advancement of hate speech detection, recent research has shifted from post-level detection to more fine-grained span-level detection. In this work, we focus on building resources for fine-grained Chinese hate speech detection. Firstly, we present the first span-level target-aware toxicity extraction dataset (STATE ToxiCN). Based on this dataset, we evaluate the performance of LLMs in span-level Chinese hate speech detection. Secondly, we provide the first Chinese hateful slang lexicon with interpretable annotations. All the hateful slang terms are sourced from real-world online platforms. Using this lexicon, we evaluate the impact of hateful slang on model detection capabilities and assess LLMs’ understanding of hateful slang. Experimental results show that current LLMs still face challenges in effectively addressing both span-level hate speech detection and the understanding of hate semantics. We hope that our resources and benchmarks will be valuable for researchers in this field.